A chunk of plastic is weighed on a scale and has a mass of 68.29 g. Next, the chunk is submerged in a graduated cylinder with an initial volume of water of 100.0 mL. After the plastic is added, the new volume in the cylinder is 129.5 mL. Find the density of the plastic in g/cm3.


1
Expert's answer
2022-04-19T00:01:04-0400

V = Vf - Vi = 129.5 mL - 100.0 mL = 29.5 mL or 29.5 cm3

"\rho" = m / V = 68.29 g / 29.5 cm3 = 2.315 g/cm3